OpenStudy (anonymous):

Determine the slope of the line through the given points (-4,1), (-1,-2). State whether the line is the graph of an increasing function, decreasing function, a constant function or not a function.

OpenStudy (anonymous):

The answer I got was: 2: increasing

OpenStudy (anonymous):

do m=(y2-y1)/(x2-x1) to find your slope then plug in one of your known points into the point slope formula to find the equation of your line

OpenStudy (anonymous):

I took m=y2-y1= -2-1=-3 x2-x1= -1-(-4) = 3 Therefore, -3/3=2:increasing

OpenStudy (anonymous):

-3/3 is -1 and therefore should be decreasing

OpenStudy (anonymous):

Okay, I got it. Thanks a lot

OpenStudy (anonymous):

np